OpenStudy (anonymous):

What is the relative maximum and minimum of the function? f(x) = 2x2 + 28x - 8 A. Minimum Value: -7 Range y > -7 B. Minimum Value: -106 Range y > -7 C. Minimum Value: 7 Range y > 7

OpenStudy (anonymous):

i think the answer is c but idk

OpenStudy (anonymous):

f(x) = 2x2 + 28x - 8 =2(x^2+14x-4) =2(x^2+14x+49-49-4) =2(x+7)^2+2(-53) =2(x+7)^2-106

OpenStudy (anonymous):

So the answer would be B?